A man save 8 Anita for 1 day 16 for 2 day 32 for 3day how many day will it take him to save the sum of 2040
Answer by Boreal(15235)

Sn=a(1-r^n)/(1-r), where 8 is 8 and r is 2
2040=8(1-2^n)/-1
255/-1=1-2^n
-2^n=-256
2^n=256
n=8 days
8/16/32/64/128/256/512/1024